OpenStudy (anonymous):

A student needs to make a rectangular cardboard piece that is 12 inches long. The area of the cardboard piece should be between 60 square inches to 300 square inches. The function f(w) relates the area of the cardboard, in square inches, to the width w in inches. Which of the following shows a reasonable domain for f(w)? A. 2 ≤ w ≤ 5 B. 5 ≤ w ≤ 25 C. 25 ≤ w ≤ 60 D. 60 ≤ w ≤ 300

OpenStudy (anonymous):

@princeharryyy @Kayla_Is_Smart @Michele_Laino @pooja195 @Ganpat

OpenStudy (michele_laino):

since w=300/12=25 and w=60/12=4 I think should be \[4\le w \le25\]

OpenStudy (ganpat):

Area of rectangle = x * y A = 60 to 300 x = 12 calculate for y.. for both the figures of area.. i.e. 60 and 300 you will find you range...

OpenStudy (anonymous):

I think D

OpenStudy (michele_laino):

oops sorry \[5\le w \le 25\]

OpenStudy (michele_laino):

@IndianBoyNiko plese note that w is the width not area